OpenStudy (anonymous):

simplify ! 1) 23/16= 2) 9/8= 3) 5/4= 4) 31/16= 5) 17/16= 6) 11/8= 7) 43/16= 8) 41/8= 9) 55/16= 10) 99/16= 11) 27/8= 12) 11/4= 13) 91/16= 14) 13/8= 15) 15/4= 16) 77/8= 17) 159/16= 19) 79/16=

OpenStudy (anonymous):

thats so much

OpenStudy (anonymous):

its only 19 lol

OpenStudy (unklerhaukus):

1) \[\frac{23}{16}=\frac{16+7}{16}=\frac{16}{16}+\frac 7{16}=1+\frac 7{16}\]

OpenStudy (unklerhaukus):

2) \[\frac{9}{8}=\frac{8+1}{8}=\frac{8}{8}+\frac{1}{8}=\quad\dots\]

OpenStudy (unklerhaukus):

3) \[\frac54=\frac{4+1}{4}=\quad\dots\quad=\quad\dots\]

OpenStudy (unklerhaukus):

... ... ...

OpenStudy (unklerhaukus):

can you see how convert improper fraction to mixed numerals ?

OpenStudy (anonymous):

If simplify means to reduce a fraction, then all of the fractions are in their reduced form. The numerators are all odd numbers and the denominators are powers of 2.

OpenStudy (anonymous):

yeah i know thanks ^^